OpenStudy (anonymous):

[3 -sqrt(5+x)] / (x-4). find limit as x reaches 4.

OpenStudy (anonymous):

the answer is -1/6, but i don't know how!

OpenStudy (anonymous):

There's the trial option, where you pick values of x that are close to 4, and see what number comes out. For example, x=3: f(x) = 0.172 x=5: f(x) = 0.162 x=3.9999: f(x) = 0.16667 x=4.0001: f(x) = 0.16667

OpenStudy (anonymous):

Oh also, L'hopital's Rule can be used, if you're studying calculus \[\lim_{x \rightarrow 4} \frac{f(x)}{g(x)} = \lim_{x \rightarrow 4} \frac{f'(x)}{g'(x)}\]

OpenStudy (anonymous):

If\[f(x) = 3 - \sqrt{5+x} \text{ then } f'(x) = \frac{1}{2\sqrt{5+x}}\]If\[g(x)=x-4 \text{ then } g'(x) = 1\]

OpenStudy (anonymous):

Sorry, f'(x) is negative.

OpenStudy (anonymous):

thanks